a) What is the final velocity of the system?

b) What is the impulse that the blue ball applies on the red ball upon impact?

c) In what direction will the impulse be?

v, 6.0 m/s v4.0 m/s m1 = 4.0 kg m2= 2.0 kg M-6.0 kg After the Collision Before the Collision
Answer #1

mass of red ball = m1 = 4 kg

                                      = v1 = 6 m/sec

mass of blue ball = m2 = 2 kg

                   v2 = - 4 m/sec

from conservation of momentum m1 v1 + m2 v2 = ( m1 + m2 ) V

                                                           4 * 6 - 2 * 4 = 6 v

                                                                   16 = 6 v

                                        final speed = V = 16 / 6 = 2.67 = 2.7 m/sec

b) impulse on red ball = change in momentum of red ball

                                 m1 V - m1 v1 = 4 * 2.7 - 4* 6 = - 13.2 kg-m/sec

c) direction is towards left direction